An aqueous KNO3 solution is made using 77.8 g of KNO3 diluted to a total solution volume of 1.98 L . (Assume a density of 1.05 g/mL for the solution.)

Calculate the molarity of the solution.

Calculate molality of the solution.

Calculate the mass percent of the solution.
Answer #1

moles of KNO3 = 77.8 g /101.1 g/mol=0.77 moles

nowmolarity equals to  number of moles of solute per liter of solution (moles/Liter)

Molarity = 0.77/ 1.98 L= 0.388 moles/lt.

mass solution = 1.05 g/mL x 1980 mL = 2079 g
mass water = 2079 - 77.8 =2001 g => 2.01 Kg

now Molality is a property of a solution and is defined as the number of moles of solute per kilogram of solvent.
molality = 0.77 mol/ 2.01=0.38m

% by mass =77.8 x 100 / 2001 =3.88 percent.

To approve a single suggestion, mouse over it and click "✔"
Click the bubble to approve all of its suggestions.
to use Ginger
Limited mode
Mass solution = 1.05 g/ml x 1980 ml
×